OpenStudy (anonymous):

f(t)=3t^2-2 find f(k-2)

OpenStudy (anonymous):

replace \(t\) by \(k-2\) be sure to use parentheses

OpenStudy (anonymous):

so if \(f(t)=3t^2-2\) then \(f(k-2)=3(k-2)^2-2\) you can clean that up with some algebra if you like

OpenStudy (anonymous):

ok so then it would be 3k^2+12-2

OpenStudy (anonymous):

no, you have to square first, multiply second, then combine like terms

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[f(k-2)=3(k-2)^2-2=3(k-2)(k-2)-2\] etc there is algebra to do

OpenStudy (anonymous):

3(k−2)^2−2 would be 3(k-2)(k-2) -2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes exactly

OpenStudy (anonymous):

3(k^2-2k-2k+4)-2 =3k^2-12k+10

OpenStudy (anonymous):

You got it :D

OpenStudy (anonymous):

are u kidding me, i got a warning for my steps?? this is rediculous